OpenStudy (anonymous):

Which graph represents the function of f(x) = the quantity of 9 x squared plus 9 x minus 18, all over 3 x plus 6? plz help me @myininaya

myininaya (myininaya):

\[f(x)=\frac{9x^2+9x-18}{3x+6}\] is what you mean?

OpenStudy (anonymous):

correct :)

myininaya (myininaya):

I suggest factoring everything

myininaya (myininaya):

the top and the bottom

myininaya (myininaya):

3x+6= what in factored form?

OpenStudy (anonymous):

idk how to do that sorry :/

myininaya (myininaya):

3x and 6 have what as a common factor?

OpenStudy (anonymous):

wat do u mean?

myininaya (myininaya):

if we write the factors of both 3x and 6 we see that 3x=3*x and 6=3*2 so 3x has factors 3 and x and 6 has factors 3 and 2 therefore the only factor tht 3x and 6 have in common is 3 therefore we can factor 3x+6 by factoring out 3 so 3x+6=3(x+2)

myininaya (myininaya):

I guess since you have never done that you will not be able to tell how to factor the numerator

OpenStudy (anonymous):

ok

myininaya (myininaya):

like you don't know how to factor \[9x^2+9x-18?\]

OpenStudy (anonymous):

i think i do

myininaya (myininaya):

oh okay let's see if you can factor that

OpenStudy (anonymous):

i'm stuck.. :/

myininaya (myininaya):

first did you factor out what all three terms had in common?

myininaya (myininaya):

9x^2=9*x^2 9x=9*x -18=9*(-2)

myininaya (myininaya):

so you can write \[9x^2+9x-18 \text{ as } 9(x^2+x-2)\]

myininaya (myininaya):

can you factor the x^2+x-2?

myininaya (myininaya):

what are two integer numbers that multiply to be -2 and add up to be 1?

OpenStudy (anonymous):

2 and -1

myininaya (myininaya):

so the factors are (x+2) and (x-1)

myininaya (myininaya):

so 9x^2+9x-18 is 9(x+2)(x-1)

OpenStudy (anonymous):

i always have trouble figuring out the numbers next to x while factoring

myininaya (myininaya):

\[f(x)=\frac{9(x+2)(x-1)}{3(x+2)}\] so we have a discontinuity at what x value?

myininaya (myininaya):

we cannot divide by 0

myininaya (myininaya):

so x+2 cannot be 0 but x+2=0 when x=?

OpenStudy (anonymous):

-2

myininaya (myininaya):

\[f(x)=\frac{9(x+2)(x-1)}{3(x+2)}=3(x-1), x \neq -2 \]

myininaya (myininaya):

our function will not exist at x=-2

myininaya (myininaya):

the x+2's canceled

myininaya (myininaya):

and 9/3=3

myininaya (myininaya):

do you think you can graph f(x)=3x-3 when x doesn't equal 2

OpenStudy (anonymous):

so in a graph its supposed to look like this?